# Language Models as Semantic Teachers: Post-Training Alignment for Medical Audio Understanding

### v1

@@ Abstract
Pre-trained audio models excel at detecting acoustic patterns in auscultation sounds but often fail to grasp their clinical significance, limiting their use and performance in diagnostic tasks. To bridge this gap, we introduce AcuLa (Audio–Clinical Understanding via Language Alignment), a lightweight post-training framework that instills semantic understanding into any audio encoder by aligning it with a medical language model, which acts as a “semantic teacher.” To enable alignment at scale, we construct a large-scale dataset by leveraging off-the-shelf large language models to translate the rich, structured metadata accompanying existing audio recordings into coherent clinical reports. Our alignment strategy combines a representation-level contrastive objective with a self-supervised modeling, ensuring that the model learns clinical semantics while preserving fine-grained temporal cues. AcuLa achieves state-of-the-art results across 18 diverse cardio-respiratory tasks from 10 different datasets, improving the mean AUROC on classification benchmarks from 0.68 to 0.79 and, on the most challenging COVID-19 cough detection task, boosting the AUROC from 0.55 to 0.89. Our work demonstrates that this audio-language alignment transforms purely acoustic models into clinically-aware diagnostic tools, establishing a novel paradigm for enhancing physiological understanding in audio-based health monitoring.
Figure 1: Performance comparison of audio-based models. (a) Average AUROC for respiratory classification tasks (T1-T9). (b) Average MAE for lung function estimation tasks (T10-T16). Our model (pink) outperforms all baselines, achieving the highest AUROC (0.79) and lowest MAE (0.82).
@@ I Introduction
Existing audio encoders capture subtle temporal and spectral patterns in auscultation sounds but still lack explicit clinical semantics. This capability gap renders them “semantically blind,” limiting their utility in high-stakes diagnostic tasks. This creates a fundamental paradox: while large language models (LLMs) possess deep, contextual knowledge of medical terminology like “systolic murmurs” or “wheezes,” this understanding remains disconnected from the audio models that process the raw signals. Digital stethoscopes and other sensors may capture a wealth of acoustic data, but without a bridge to semantic meaning, this information remains underutilized.
@@
Multimodal contrastive learning, popularized by frameworks like CLIP [radford2021learning], aims to bridge such divides by aligning heterogeneous modalities in a shared embedding space. This approach has enabled powerful cross-modal retrieval and classification capabilities. However, these methods often suffer from a persistent ’modality gap,’ where embeddings from different sources form distinct clusters, hindering fine-grained alignment and interpretability [liang2022mind, ray2023cola]. This issue is particularly acute in clinical applications, where subtle acoustic variations carry significant diagnostic weight and demand precise semantic grounding.
@@
While existing methods have sought to bridge this gap through architectural modifications [chen2023revisiting], auxiliary objectives [lee2022uniclip], or post-training alignment [yamaguchi2025post], their focus has been on aligning two perceptual modalities. Even recent advances in knowledge transfer, such as [gan2025cross], operate under this paradigm, enhancing language models by injecting knowledge from vision models. These approaches share a common directional assumption: knowledge flows from concrete perception to abstract representation. Our work fundamentally diverges by framing the problem not as mutual alignment, but as a directed and asymmetric knowledge infusion. We investigate how to leverage the vast, pre-existing semantic knowledge of an LLM, treating it as a “semantic teacher” to guide and enrich a specialized “acoustic student.” This presents a distinct challenge: grounding high-level clinical concepts from text into the fine-grained, temporal patterns of a raw audio signal, a problem unexplored by prior methods.
@@
This frontier is especially important for domains rich with temporal and semantic information, such as medical audio. Millisecond-scale events like the onset of a lung crackle or the specific timing of a heart murmur contain precise clinical information that current audio-only models struggle to link to a diagnosis. To address this, we introduce AcuLa (Audio–Clinical Understanding via Language Alignment), a general, post-training framework that instills clinical semantic understanding into any pre-trained audio encoder. In our approach, a frozen, LLM acts as a ’semantic teacher,’ guiding an audio ’student’ model to map acoustic patterns to their corresponding clinical meanings.
@@
We demonstrate AcuLa’s effectiveness in the challenging domain of cardio-respiratory health (see Figure 1). By synthetically generating a large-scale dataset of clinical reports from structured metadata, we create the necessary paired data to align audio recordings with their clinical interpretations. Our results show that this alignment transforms standard audio encoders into clinically-aware models that can better differentiate subtle pathologies and significantly improves downstream tasks performance.
@@ II-A Medical Audio Analysis
The analysis of respiratory and cardiac sounds has traditionally been a unimodal task. Supervised deep learning models have shown success in disease classification given sufficient expert-labeled data, while self-supervised pre-training on large, unannotated corpora has proven effective in low-resource settings [zhang2024towards]. Recognizing the limitations of audio-only analysis, recent efforts have started to incorporate textual features. For instance, RespLLM [zhang2024respllm] fuses spectral features with clinical notes via cross-modal attention. However, such approaches typically treat audio and text as separate processing streams, fusing them only at a late stage for a final decision. This shallow integration prevents the audio encoder itself from learning a semantically rich representation, limiting its ability to ground acoustic events in clinical meaning. Our work moves beyond late fusion by directly aligning audio and language representations at a deep feature level.
@@ II-B Cross-Modal Alignment
Representation Alignment. The dominant paradigm, popularized by CLIP [radford2021learning], learns a shared embedding space where corresponding pairs from different modalities are projected to be close. This contrastive approach has been successfully extended to the audio domain with models like CLAP [elizalde2023clap] and AudioCLIP [guzhov2022audioclip], enabling better cross-modal learning. However, these frameworks often struggle with a “modality gap,” where representations remain clustered by their original modality, hindering fine-grained understanding [liang2022mind]. This limitation is particularly critical for medical signals, where subtle pattern differences are diagnostically vital.
Knowledge Transfer and Distillation. An alternative perspective is to view alignment as a form of directed knowledge transfer. This can be achieved through various mechanisms. Generative methods like AudioLM [borsos2023audiolm] and AudioGen [kreuk2022audiogen] learn to synthesize audio conditioned on text, implicitly learning a shared structure. Knowledge distillation [hinton2015distilling] transfers representations from a “teacher” to a “student” network mostly focusing on supervised task-specific models. Most relevant to our work is regularization-based alignment, where a model’s representation space is regularized to match that of another. CMAR [gan2025cross], for example, regularizes an LLM’s latent space using features from a vision model. Recent connector-based approaches for speech–text learning [tan-etal-2025-ssr, yu-etal-2023-speech] instead focus on token-level correspondences, training cross-modal adapters or attention modules end to end and often updating the language model itself. These methods require paired audio–text data and emphasize fine-grained synchronous alignment. In contrast, our setting targets post-training global embedding alignment: the medical LLM remains frozen, and lightweight projection heads align encoder representations under schema-constrained, clip-level textual supervision.
Our Contributions and Positioning. Despite these advances, two critical gaps remain. First, within medical audio, a deep, semantic alignment between audio representations and clinical text is underexplored. Second, in the broader field of cross-modal learning, knowledge transfer has been almost exclusively studied from perceptual domains, like vision to language, or between static modalities like images and text, with few attempts at unifying audio like ImageBind [girdhar2023imagebind]. AcuLa is the first work to bridge both gaps. We introduce a framework for deep representation alignment specifically for temporal medical audio. Crucially, we invert the conventional direction of knowledge transfer, using a pre-trained LLM as a semantic teacher to instill clinical understanding into a specialized audio encoder, directly addressing the challenges of grounding abstract semantics in time-varying signals.
@@ III-C Training Objective
Here, we optimize the audio encoder parameters $\theta$ and the projection head parameters $\psi$, while the LLM parameters $\phi$ remain frozen.
Semantic Alignment via Centered Kernel Alignment (CKA). To align the two modalities, we require a robust similarity metric between the sets of batch embeddings, $\mathbf{H}^{\text{audio}}=[\mathbf{h}^{\text{audio}}_{1},\dots,\mathbf{h}^{\text{audio}}_{B}]^{T}$ and $\mathbf{H}^{\text{language}}=[\mathbf{h}^{\text{language}}_{1},\dots,\mathbf{h}^{\text{language}}_{B}]^{T}$. We use Centered Kernel Alignment (CKA) [kornblith2019similarity], a metric that compares the geometric structure of representation spaces, making it invariant to isotropic scaling and rotation. CKA is defined via the Gram matrices of the mean-centered representations, $\mathbf{G}^{\text{audio}}=\bar{\mathbf{H}}^{\text{audio}T}\bar{\mathbf{H}}^{\text{audio}}$ and $\mathbf{G}^{\text{language}}=\bar{\mathbf{H}}^{\text{language}T}\bar{\mathbf{H}}^{\text{language}}$:
@@
Acoustic Preservation via Self-Supervised Modeling (SSM). The alignment loss alone could cause the audio encoder to discard acoustic information not present in the simplified text reports, a phenomenon known as representation collapse [jing2021understanding]. To counteract this and preserve fine-grained temporal details, we incorporate the audio encoder’s self-supervised objective, $\mathcal{L}_{\text{SSM}}$. For many state-of-the-art audio models, this is a masked acoustic modeling loss [huang2022masked], which involves reconstructing masked patches of the input spectrogram or contrastive objective [saeed2021contrastive]. In our case, they act as a powerful regularizer, ensuring the audio representations do not catastrophically forget their core acoustic modeling capabilities. The final optimization thus becomes:
@@ III-D Synthetic Alignment Data Generation
A significant challenge in medical multimodal learning is the scarcity of large-scale, paired audio-text datasets. To overcome this, we devise a scalable strategy to generate high-quality clinical text from existing audio datasets with structured metadata. We leverage an off-the-shelf LLM, GPT-4o [openai2024gpt4technicalreport], to synthesize clinical reports. For each audio recording from public datasets like ICBHI [DVN/HT6PKI_2023] and Circor [oliveira2021circor], we compile its available metadata—patient demographics, recording conditions, and diagnostic labels (e.g., presence of crackles/wheezes)—into a structured prompt (see Appendix A and Figure 3). The LLM is tasked to act as a clinical specialist (e.g., a pulmonologist) and generate a concise, natural-language report based only on the provided information. With precise and explicit prompting of LLM, we ensure factual grounding while encouraging linguistic diversity. Examples of metadata–report pairs are provided in Appendix B. This process yields a corpus of over $100,000$ paired audio-report samples, summarized in Table I.
@@
TABLE II: Downstream task characteristics grouped by task category. Abbreviations: Exhal. = Exhalation, Obstr. = Obstructive, Resp. = Respiratory, Sam. = Samples, Sub. = Subjects.
@@ IV-A Baselines
To situate AcuLa’s performance, we compare it against a diverse set of strong pre-trained models representing different architectural and training paradigms. These include VGGish, AudioMAE [huang2022masked], and CLAP [elizalde2023clap]. We also include the OPERA family of models (generative and contrastive both) [zhang2024towards], which are foundation models trained specifically on respiratory audio. As a non-deep-learning benchmark, we use OpenSMILE [eyben2010opensmile] to extract a standard set of hand-crafted acoustic features. For all deep learning baselines, we use the authors’ official pre-trained encoders to extract features. Unless otherwise specified, AcuLa is applied to the OPERA (GT) encoder to demonstrate its enhancement capabilities.
TABLE III: AUROC ($\uparrow$) on health condition inference tasks (higher is better). The best model for each task is highlighted. We report mean and standard deviation from five independent runs. All baseline results (VGGish, AudioMAE, CLAP, OCT, OCE, OGT) are reported in [zhang2024towards]. $\checkmark$ denotes when our method outperforms the OpenSmile baseline (detailed in Appendix C), while * indicates superior performance compared to all other pretrained models.
@@
TABLE IV: MAE ($\downarrow$) on lung function estimation tasks (lower is better). Best model per task is highlighted. We report mean and standard deviation across subjects. All baseline results (VGGish, AudioMAE, CLAP, OCT, OCE, OGT) are from [zhang2024towards]. $\checkmark$ denotes when our method outperforms the OpenSmile baseline (detailed in Appendix C), while * indicates superior performance compared to all other pretrained models.
@@ IV-B Implementation Details
Our implementation of AcuLa employs MedGemma-4B [google_medgemma_2025] as the default language model ($\mathcal{L}_{\phi}$), which has been pre-trained on medical literature, and utilizes the OPERA encoder [zhang2024towards] as the audio foundation model ($\mathcal{A}_{\theta}$). To bridge the modality gap, we introduce two MLP projection heads. The audio projection MLP maps the 384-dimensional OPERA features to a 512-dimensional shared space via a two-layer network (384 $\rightarrow$ 1024 $\rightarrow$ 512) with ReLU activation and 20% dropout. The language projection MLP similarly transforms the 2048-dimensional MedGemma-4B hidden states to the same 512-dimensional space.
During the alignment phase, audio inputs are preprocessed into 8-second segments sampled at 16kHz and converted to log-mel spectrograms with 64 mel bins. We apply on-the-fly data augmentation using the AugLy [papakipos2022augly] library, randomly selecting one transformation per sample from a set including a 5dB volume increase, amplitude normalization, low-pass filtering (300Hz cutoff), or high-pass filtering (3000Hz cutoff). We train the model using the AdamW optimizer, applying a learning rate of $1\times 10^{-5}$ to both the audio encoder and the projection heads. We train for 50 epochs with a linear learning rate schedule incorporating 400 warmup steps. The batch size is set to 24, with gradient accumulation over 2 steps to fit within memory constraints. The combined loss function weights the CKA-based alignment loss and the SSM loss equally ($\lambda_{\text{align}}=\lambda_{\text{MAM}}=1.0$). The entire alignment process takes approximately 30 hours on a single NVIDIA A100 GPU.

### v2

@@ Abstract
Pre-trained audio models excel at detecting acoustic patterns in auscultation sounds but often fail to grasp their clinical significance, limiting their use and performance in diagnostic tasks. To bridge this gap, we introduce AcuLa (Audio–Clinical Understanding via Language Alignment), a lightweight post-training framework that instills semantic understanding into any audio encoder by aligning it with a medical language model, which acts as a “semantic teacher.” To enable alignment at scale, we construct a large-scale dataset by leveraging off-the-shelf large language models to translate the rich, structured metadata accompanying existing audio recordings into coherent clinical reports. Our alignment strategy combines a representation-level contrastive objective with a self-supervised modeling, ensuring that the model learns clinical semantics while preserving fine-grained temporal cues. AcuLa achieves state-of-the-art results across 18 diverse cardio-respiratory tasks from 10 different datasets, improving the mean AUROC on classification benchmarks from 0.68 to 0.79 and, on the most challenging COVID-19 cough detection task, boosting the AUROC from 0.55 to 0.89. Our work demonstrates that this audio-language alignment transforms purely acoustic models into clinically-aware diagnostic tools, establishing a novel paradigm for establishing a paradigm for injecting clinical-language semantics into audio representations for audio-based health monitoring.
@@ I Introduction
Existing audio encoders capture subtle temporal and spectral patterns in auscultation sounds but still lack explicit clinical semantics. This leaves them “semantically blind,” limiting their use in high-stakes diagnostic tasks. A fundamental paradox follows: while large language models (LLMs) understand medical concepts such as “systolic murmurs” and “wheezes,” this knowledge remains disconnected from the audio models that process raw signals. Digital stethoscopes and other sensors can capture rich acoustic data, but without a bridge to semantic meaning, much of this information remains underused.
@@
Multimodal contrastive learning, popularized by frameworks such as CLIP [24], seeks to bridge such divides by aligning heterogeneous modalities in a shared embedding space. This has enabled strong cross-modal retrieval and classification, but these methods often suffer from a persistent “modality gap,” where embeddings from different modalities form distinct clusters, limiting fine-grained alignment and interpretability [20, 25]. This issue is especially critical in clinical settings, where subtle acoustic differences can carry major diagnostic significance and require precise semantic grounding.
@@
Existing approaches have addressed this gap through architectural changes [3], auxiliary objectives [19], or post-training alignment [30], but they mainly focus on aligning two perceptual modalities. Even recent knowledge transfer methods such as [8] follow this paradigm, improving language models with knowledge from vision models. These approaches assume that knowledge flows from concrete perception to abstract representation. In contrast, our work frames the problem as directed, asymmetric knowledge infusion. We leverage the broad semantic knowledge of an LLM as a “semantic teacher” to guide and enrich a specialized “acoustic student.” This introduces a distinct challenge: grounding high-level clinical concepts from text into the fine-grained temporal patterns of raw audio, which remains largely unexplored.
@@
This frontier is especially important for domains rich with temporal and semantic information, such as medical audio. Millisecond-scale events like the onset of a lung crackle or the specific timing of a heart murmur contain precise clinical information that current audio-only models struggle to link to a diagnosis. To address this, we introduce AcuLa (Audio–Clinical Understanding via Language Alignment), a general, post-training framework that instills clinical semantic understanding into any pre-trained audio encoder. In our approach, a frozen, LLM acts as a “semantic teacher,” guiding an audio “student” model to map acoustic patterns to their corresponding clinical meanings.
@@
We demonstrate AcuLa’s effectiveness in the challenging domain of cardio-respiratory health. By synthetically generating a large-scale dataset of clinical reports from structured metadata, we create the necessary paired data to align audio recordings with their clinical interpretations. Our results show that this alignment transforms standard audio encoders into clinically-aware models that can better differentiate subtle pathologies and significantly improves downstream tasks performance.
@@ II-A Medical Audio Analysis
Respiratory and cardiac sound analysis has traditionally been treated as a unimodal task. Supervised deep learning performs well with sufficient expert labels, while self-supervised pre-training helps in low-resource settings [32]. More recent work has begun incorporating text; for example, RespLLM [33] combines spectral features with clinical notes through cross-modal attention. However, these methods usually process audio and text separately and fuse them only at a late stage for prediction. This limits the audio encoder’s ability to learn semantically rich representations. In contrast, our work aligns audio and language at the feature level to directly ground acoustic events in clinical meaning.
@@ II-B Cross-Modal Alignment
Representation Alignment. The dominant paradigm, popularized by CLIP [24], learns a shared embedding space where corresponding pairs from different modalities are projected to be close. This contrastive approach has been successfully extended to the audio domain with models like CLAP [5] and AudioCLIP [12], enabling better cross-modal learning. However, these frameworks often struggle with a “modality gap,” where representations remain clustered by their original modality, hindering fine-grained understanding [20]. This limitation is particularly critical for medical signals, where subtle pattern differences are diagnostically vital.
Knowledge Transfer and Distillation. Alignment can also be viewed as directed knowledge transfer. Generative methods such as AudioLM [2] and AudioGen [17] learn audio conditioned on text, implicitly inducing shared structure. Knowledge distillation [13] transfers representations from a “teacher” to a “student,” typically in supervised, task-specific settings. Most related to our work is regularization-based alignment, where one representation space is regularized to match another (e.g., CMAR [8] regularizes an LLM using features from a vision model). Connector-based speech–text approaches [28, 31] instead model token-level correspondences by training cross-modal adapters end to end and often updating the language model. In contrast, we perform post-training global embedding alignment: the medical LLM is frozen, and lightweight projection heads align audio encoder representations using schema-constrained, clip-level textual supervision.
Our Contributions and Positioning. Despite these advances, two gaps remain. First, semantic alignment between medical audio representations and clinical text is underexplored. Second, cross-modal knowledge transfer has mostly focused on static modalities (e.g., vision–language), with limited attempts to unify audio (e.g., ImageBind [9]). AcuLa bridges these gaps by introducing a lightweight framework for representation alignment tailored to temporal medical audio, using a pre-trained LLM as a semantic teacher for adapting a specialized audio encoder. The frozen LLM provides clinical-language semantics as a regularizer; we do not claim physiological ground truth, causal understanding, or clinical decision-making.
@@ III-C Training Objective
Here, we optimize the audio encoder parameters $\theta$ and the projection head parameters $\psi$, while the LLM parameters $\phi$ remain frozen. Unless stated otherwise, we set $\lambda_{\text{align}}=\lambda_{\text{SSM}}=1.0$.
Semantic Alignment via Centered Kernel Alignment (CKA). To align the two modalities, we require a robust similarity metric between the sets of batch embeddings, $\mathbf{H}^{\text{audio}}=[\mathbf{h}^{\text{audio}}_{1},\dots,\mathbf{h}^{\text{audio}}_{B}]^{T}$ and $\mathbf{H}^{\text{language}}=[\mathbf{h}^{\text{language}}_{1},\dots,\mathbf{h}^{\text{language}}_{B}]^{T}$. We obtain $\mathbf{h}^{\text{language}}$ by mean-pooling the last-layer hidden states of the frozen LLM, and map both modalities to a shared embedding space via two MLP projectors. We use Centered Kernel Alignment (CKA) [16], a metric that compares the geometric structure of representation spaces, making it invariant to isotropic scaling and rotation. CKA is defined via the Gram matrices of the mean-centered representations, $\mathbf{G}^{\text{audio}}=\bar{\mathbf{H}}^{\text{audio}T}\bar{\mathbf{H}}^{\text{audio}}$ and $\mathbf{G}^{\text{language}}=\bar{\mathbf{H}}^{\text{language}T}\bar{\mathbf{H}}^{\text{language}}$:
@@
We use the LLM as a source of clinical-language semantics for regularization and do not interpret this alignment as physiological ground truth or causal biomedical understanding. The alignment is performed on clip-level global embeddings and does not explicitly model event-level temporal–semantic correspondence within a recording.
Acoustic Preservation via Self-Supervised Modeling (SSM). The alignment loss alone may cause the audio encoder to discard acoustic information not captured by simplified text reports (representation collapse [15]). We therefore include the audio encoder’s self-supervised objective, $\mathcal{L}_{\text{SSM}}$. In our implementation, $L_{SSM}$ is instantiated as the audio encoder’s masked acoustic reconstruction loss. We use the encoder’s default masked reconstruction setup: the mel-spectrogram is split into $4\times 4$ patches, $70\%$ are masked uniformly at random. For many state-of-the-art audio models, this is a masked acoustic modeling loss [14] (or a contrastive objective [26]), acting as a regularizer to preserve acoustic modeling capability. The final optimization thus becomes:
@@ III-D Synthetic Alignment Data Generation
A significant challenge in medical multimodal learning is the scarcity of large-scale, paired audio-text datasets. To overcome this, we devise a scalable strategy to generate high-quality clinical text from existing audio datasets with structured metadata. We leverage an off-the-shelf LLM, GPT-4o [22], to synthesize clinical reports. For each audio recording from public datasets like ICBHI [27] and Circor [21], we compile its available metadata—patient demographics, recording conditions, and diagnostic labels (e.g., presence of crackles/wheezes)—into a structured prompt (see Appendix A and Figure 2). The LLM is tasked to act as a clinical specialist (e.g., a pulmonologist) and generate a concise, natural-language report based only on the provided information. With precise and explicit prompting of LLM, we ensure factual grounding while encouraging linguistic diversity. Examples of metadata–report pairs are provided in Appendix B. This process yields a corpus of over $100,000$ paired audio-report samples, summarized in Table I.
@@
III-E Expert Verification of Synthetic Reports
To validate the LLM-generated synthetic reports used for alignment, we conducted an expert review of 50 metadata–report pairs sampled from the synthesized corpus (10 from each of 5 datasets), covering diverse respiratory and cardiac scenarios. Each pair included the structured metadata and its generated clinical report. A pulmonologist evaluated metadata consistency, unsupported added detail, and clinical plausibility. We define accuracy as the proportion of reports fully consistent with the source metadata, and unsupported-detail rate as the proportion containing information not directly supported by the metadata. Results are summarized in Table II.
As shown in Table II, all 50 reviewed reports were fully consistent with the source metadata and clinically plausible. Unsupported added detail appeared in 6 cases (12%), but these did not contradict the metadata and mainly reflected mild over-specification. For example, one clinician noted that the phrase “no abnormalities in respiratory function” overstated what local normal breath sounds alone can support. Overall, the results show that the generated reports provide high-fidelity, clinically plausible metadata-grounded semantic supervision for alignment.
TABLE II: Expert verification results for LLM-generated synthetic clinical reports. Accuracy is defined as metadata consistency, and unsupported-detail rate denotes the proportion of reports containing information not directly supported by the structured input metadata.
Metric Count Percentage
Metadata consistency (Accuracy) 50/50 100%
No unsupported added detail 44/50 88%
Unsupported-detail rate 6/50 12%
Clinical plausibility 50/50 100%
TABLE III: Downstream task characteristics grouped by task category. Abbreviations: Exhal. = Exhalation, Obstr. = Obstructive, Resp. = Respiratory, Sam. = Samples, Sub. = Subjects.
@@ IV-A Baselines
To situate AcuLa’s performance, we compare it against a diverse set of strong pre-trained models representing different architectural and training paradigms. These include VGGish, AudioMAE [14], and CLAP [5]. We also include the OPERA family of models (generative and contrastive both) [32], which are foundation models trained specifically on respiratory audio. As a non-deep-learning benchmark, we use OpenSMILE [7] to extract a standard set of hand-crafted acoustic features. For all deep learning baselines, we use the authors’ official pre-trained encoders to extract features. Unless otherwise specified, AcuLa is applied to the OPERA (GT) encoder to demonstrate its enhancement capabilities.
TABLE IV: AUROC ($\uparrow$) on health condition inference tasks (higher is better). The best model for each task is highlighted. We report mean and standard deviation from five independent runs. All baseline results (VGGish, AudioMAE, CLAP, OCT, OCE, OGT) are reported in [32]. $\checkmark$ denotes when our method outperforms the OpenSmile baseline (detailed in Appendix C), while * indicates superior performance compared to all other pretrained models.
@@
TABLE V: MAE ($\downarrow$) on lung function estimation tasks (lower is better). Best model per task is highlighted. We report mean and standard deviation across subjects. All baseline results (VGGish, AudioMAE, CLAP, OCT, OCE, OGT) are from [32]. $\checkmark$ denotes when our method outperforms the OpenSmile baseline (detailed in Appendix C), while * indicates superior performance compared to all other pretrained models.
@@ IV-B Implementation Details
Our implementation of AcuLa employs MedGemma-4B [10] as the default language model ($\mathcal{L}_{\phi}$), which has been pre-trained on medical literature, and utilizes the OPERA encoder [32] as the audio foundation model ($\mathcal{A}_{\theta}$). To bridge the modality gap, we introduce two MLP projection heads. The audio projection MLP maps the 384-dimensional OPERA features to a 512-dimensional shared space via a two-layer network (384 $\rightarrow$ 1024 $\rightarrow$ 512) with ReLU activation and 20% dropout. The language projection MLP similarly transforms the 2048-dimensional MedGemma-4B hidden states to the same 512-dimensional space.
During the alignment phase, audio inputs are preprocessed into 8-second segments sampled at 16kHz and converted to log-mel spectrograms with 64 mel bins. We apply on-the-fly data augmentation using the AugLy [23] library, randomly selecting one transformation per sample from a set including a 5dB volume increase, amplitude normalization, low-pass filtering (300Hz cutoff), or high-pass filtering (3000Hz cutoff). We train the model using the AdamW optimizer, applying a learning rate of $1\times 10^{-5}$ to both the audio encoder and the projection heads. We train for 50 epochs with a linear learning rate schedule incorporating 400 warmup steps. The batch size is set to 24, with gradient accumulation over 2 steps to fit within memory constraints. The combined loss function weights the CKA-based alignment loss and the self-supervised modeling loss equally ($\lambda_{align}=\lambda_{SSM}=1.0$). In our implementation, $L_{SSM}$ is instantiated as the masked acoustic reconstruction objective of the audio encoder.The entire alignment process takes approximately 30 hours on a single NVIDIA A100 GPU.